OpenStudy (anonymous):

sam will place 3 of his 8 modeled trucks on a self how many diffrent orders can he place 3 of them on the self

OpenStudy (dumbcow):

permutations 8P3 = 8!/(8-3)! = 8!/5! = 8*7*6 = 336